OpenStudy (anonymous):

-7/10 divided by 2/5? I dont want just an answer, id you can explain that would be awesome!

OpenStudy (anonymous):

You would first need to multiply 2 and 5 by 2. Because 5 x2=10 and you need a common denominator. That would be 4/10. With fractions, you multiply when trying to divide. you multiply using the reciprical which is just the fraction flip flopped. This would give you -7 x 4 since you multiply just the top and keep the denominator (10) the same. -7 x 4= -28/10 or -2 8/10 :) If you want to simplify it's -2 and 4/5

OpenStudy (anonymous):

-7/10 = -.7 2/5=.4 -.7/.4=-1.75